$SPAD/src/input series3.input

Nicolas Francois and Martin Rubey

December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)spool series3.output
)set message test on
)set message auto off
)clear all

———-

Question by Nicolas Francois <nicolas.francois@free.fr>: Is there any way to obtain the decomposition in simple elements of a fraction of the form:

\frac{1}{(1 − x)(1 − x^{2})(1 − x^{5})}

Martin Rubey <martin.rubey@math.uni-hannover.de> replied:

— * —

--S 1 of 4
SAEs5 := SAE(FRAC INT, UP(s5,FRAC INT),s5^2-5)
--E 1
--S 2 of 4
p:UP(x,SAEs5) := (x^5-1)*(x^2-1)*(x-1)
--E 2
--S 3 of 4
factor p
--E 3
--S 4 of 4
partialFraction(1/p,x)
--E 4
)spool
)lisp (bye)

———-

References^{[1] nothing}

4